OpenStudy (anonymous):

Help!!! Solve for x: 3(x + 1) = -2(x - 1) + 6. 1 4 5 25

OpenStudy (anonymous):

3(x+1)=-2(x-1)+6 3x+3=-2x+2+6 3x+3=-2x+8 -3 -3 3x=-2x+5 +2x +2x 5x=5 divide both sides by 5 x=1

OpenStudy (anonymous):

thanks